# Group Invariant Quantum Latin Squares

### v1

@@ 1 Introduction
In [15], Roberson and Schmidt introduced a highly symmetric type of strategy for the graph isomorphism game. These so-called group invariant strategies are invariant under the regular action of a given group $G$ on the players’ inputs and outputs (this is made precise in Section 2). For a strategy arising from a quantum Latin square, this amounts to an invariance of the square of the modulus of the inner products of its entries under independent regular actions on its row and column indices. Currently, a complete characterization of such quantum Latin squares seems out of reach. Instead we strengthen this invariance condition to apply to the inner products themselves, rather than the squares of their moduli, and say that such quantum Latin squares are group invariant. We also consider the case where the row and column indices are acted on by different groups, say $G$ and $G^{\prime}$ respectively, and refer such Latin squares as being $(G,G^{\prime})$-invariant.
In general, there are $n^{4}$ inner products among the $n^{2}$ vectors in an $n\times n$ quantum Latin square. However, due to the symmetry of group invariant quantum Latin squares, they have at most $n^{2}$ distinct inner products, and these can be stored in an $n\times n$ matrix that we term the transformation matrix. The transformation matrix completely determines a $(G,G^{\prime})$-invariant quantum Latin square up to a global isometry, and we are able to completely characterize these matrices in terms of simple conditions based on the groups $G$ and $G^{\prime}$ (see Theorem 4.5). . This allows us to prove our main result, that the maps consisting of conjugating by a transformation matrix are precisely the trace-preserving isomorphisms of the left regular representations of $G$ and $G^{\prime}$ that commute with conjugate transpose. This implies that a $(G,G^{\prime})$-invariant quantum Latin square exists if and only if the multiset of degrees of irreducible representations of $G$ is the same as that of $G^{\prime}$, i.e., if and only if the group algebras of $G$ and $G^{\prime}$ are isomorphic. In particular, if $G$ and $G^{\prime}$ are abelian groups of the same order, such a quantum Latin square always exists and in this case there are finitely many. If, however, a $(G,G^{\prime})$-invariant quantum Latin square exists for non-abelian groups $G$ and $G^{\prime}$, then there are uncountably many of them.
@@
In Section 6 we delve into some representation theory that we will need for our main result and also investigate the relationship between transformation matrices of $(G,G^{\prime})$-invariant quantum Latin squares and what we refer to as quasi-regular representations of $G$ and $G^{\prime}$. We prove our main result in Section 7, namely that maps given by conjugating by a transformation matrix of a $(G,G^{\prime})$-invariant quantum Latin square are precisely the trace and conjugate transpose preserving isomorphisms of the group algebras of $G$ and $G^{\prime}$. Section 8 shows how to actually construct $(G,G^{\prime})$-transformation matrices, assuming one has unitaries that block-diagonalize the group algebras. In the abelian case this is achieved by the character tables of the groups, and thus we can construct all $(G,G^{\prime})$-transformation matrices since there are finitely many in this case. We study the abelian case further in Section 9 and show that in this case the convex hull of the correlations produced by $(G,G^{\prime})$-invariant quantum Latin squares is the image of the set of classical $(G,G^{\prime})$-invariant correlations under a particular unitary map. In Section 10 we show that the submatrices of $(G,G^{\prime})$-transformation matrices corresponding to the rows/columns with a single nonzero entry give isomorphisms between subgroups of $G$ and $G^{\prime}$, as well as the linear characters of these groups. Additionally, we prove a partial converse assuming a certain product structure of the groups $G$ and $G^{\prime}$. Section 11 presents some computational results and important examples, e.g., an example of a non-classical correlation produced by a $\mathbb{Z}_{2}^{4}$-invariant quantum Latin square, answering a question of [15]. Finally, we end with a discussion of our results and some open problems.
@@ 2 The Isomorphism Game
It follows from the proof of Theorem 5.4 in [1] that a quantum permutation matrix $\mathcal{P}=(p_{xy})_{x\in V(X),y\in V(Y)}$ satisfying the conditions of the above theorem results in a correlation $p$ given by
@@
where $\operatorname{tr}$ denotes the normalized trace, i.e., $\operatorname{tr}(M)=\operatorname{Tr}(M)/d$ for $M\in M_{d}(\mathbb{C})$. It is well known that if all of the entries of a quantum permutation matrix pariwise commute, then the resulting correlation is classical. The fact that the converse is not true is the focus of [15].
@@
It is likely, but unknown, that not every winning quantum correlation for the $(X,Y)$-isomorphism game arises from a quantum permutation matrix via Equation (2). However, it follows from known results in the literature, see e.g. [8], that every winning quantum correlation for the $(X,Y)$-isomorphism game is a convex combination of correlations of the form (2). Importantly, the sets of classical and quantum correlations are convex with the latter containing the former.
Notably, it is possible for graphs $X$ and $Y$ to be quantum isomorphic but not isomorphic [1]. In this case, the correlation of any winning quantum strategy for the $(X,Y)$-isomorphism game is necessarily non-classical. However, the converse does not hold: there exist non-classical quantum correlations that win the $(X,Y)$-isomorphism game, but only for isomorphic $X$ and $Y$, e.g., the correlation used in [15, Theorem 3.19] to show that $K_{5}$ has “nonlocal symmetry”. Note that if $p$ is a winning correlation for the $(X,Y)$-isomorphism game, then it is also a winning correlation for the isomorphism game where $X$ and $Y$ are replaced by empty graphs on the same vertex set. Since we are mainly focused on group invariant quantum Latin squares and the correlations they produce, this least restrictive isomorphism game is the one we will consider. Since for empty graphs, two vertices can only be equal or distinct non-adjacent, in this case it makes sense to simply think of the isomorphism game as a bijection game between two sets, in which the players must give the same answer if and only if they are given the same input. For us, the two sets will always be two groups $G$ and $G^{\prime}$. As with the isomorphism game, we restrict to the case where $|G|=|G^{\prime}|$, but remark that this assumption can be removed as in the isomorphism game case by passing to a larger game. It is then clear that the winning deterministic strategies for the $(G,G^{\prime})$-bijection game correspond precisely to the bijections between $G$ and $G^{\prime}$. Thus there is always a winning classical strategy, but we are interested in the winning quantum strategies that produce non-classical correlations. However, in Section 3.2 we will show how to take any winning group invariant correlation for the bijection game and determine all the graph isomorphism games for which it is also a winning strategy.
@@ 3.2 Cayley (Di)Graphs
For any $n\times n$ doubly stochastic matrix $D$ and vectors $u,v\in\mathbb{C}^{n}$ with the same multiset of entries, it is well known that $Du=v$ if and only if $D_{ij}=0$ whenever $u_{j}\neq v_{i}$ (see, e.g., [9, Lemma 7.1]). Therefore $D^{p}\mathbf{e}_{C^{\prime}}=\mathbf{e}_{C}$ if and only if $D^{p}_{a,b}=0$ whenever ($a\in C$ and $b\notin C^{\prime}$) or ($a\notin C$ and $b\in C^{\prime}$). By definition of the graph $B$, this is equivalent to there being no edges between $C\cup C^{\prime}$ and $(G\cup G^{\prime})\setminus(C\cup C^{\prime})$. This is, in turn, equivalent to $C\cup C^{\prime}$ being the vertex set of a union of connected components of $B$, which is equivalent to $C=C_{I}$ and $C^{\prime}=C^{\prime}_{I}$ for some $I\subseteq[k]$. ∎
@@ 7 Existence of transformation matrices
By the Wedderburn-Artin Theorem [3, Theorem 1.3.5], any semisimple Artinian algebra can be written as a direct sum of full matrix algebras whose dimensions are uniquely determined. Now the result follows from Lemma 7.1. ∎
@@ 9 The abelian case
It is straightforward to check that the transformation matrix of the $(G,G^{\prime})$-invariant quantum Latin square described in Example 4.4 is precisely the matrix $\mathcal{C}^{\dagger}P^{\pi}\mathcal{C}^{\prime}$ of the above theorem, where $\pi$ is the bijection taking $\chi^{\prime}_{i}$ to $\chi_{i}$. Thus the construction described in that example, which is the generalization of the construction of [15, Definition 4.1] to the case of possibly different groups $G$ and $G^{\prime}$, produces all $(G,G^{\prime})$-invariant quantum Latin squares, up to isometry.
@@ 10 Isomorphisms
Now consider the groups $K$ and $K^{\prime}$. It is well known that the irreducible characters of a direct product of groups are precisely the Kronecker products of the irreducible characters of the factors (see for example [6, Theorem 19.18]). It follows that the multisets of degrees of the irreducible representations of $K$ and $K^{\prime}$ are the same. Therefore, there exists a $(K,K^{\prime})$-transformation matrix, $W$. We will show that that the matrix $U:=V\otimes W$ is a $(G,G^{\prime})$-transformation matrix satisfying the required conditions. Clearly, $U\in\mathbb{C}^{(H\times K)\times(H^{\prime}\times K^{\prime})}=\mathbb{C}^{G\times G^{\prime}}$, and it is a unitary matrix since $V$ and $W$ are unitary.

### v2

@@ 1 Introduction
In [18], Roberson and Schmidt introduced a highly symmetric type of strategy for the graph isomorphism game. These so-called group invariant strategies are invariant under the regular action of a given group $G$ on the players’ inputs and outputs (this is made precise in Section 2). For a strategy arising from a quantum Latin square, this amounts to an invariance of the square of the modulus of the inner products of its entries under independent regular actions on its row and column indices. Currently, a complete characterization of such quantum Latin squares seems out of reach. Instead we strengthen this invariance condition to apply to the inner products themselves, rather than the squares of their moduli, and say that such quantum Latin squares are group invariant. We also consider the case where the row and column indices are acted on by different groups, say $G$ and $G^{\prime}$ respectively, and refer to such Latin squares as being $(G,G^{\prime})$-invariant.
In general, there are $n^{4}$ inner products among the $n^{2}$ vectors in an $n\times n$ quantum Latin square. However, due to the symmetry of group invariant quantum Latin squares, they have at most $n^{2}$ distinct inner products, and these can be stored in an $n\times n$ matrix that we term the transformation matrix. The transformation matrix completely determines a $(G,G^{\prime})$-invariant quantum Latin square up to a global isometry, and we are able to completely characterize these matrices in terms of simple conditions based on the groups $G$ and $G^{\prime}$ (see Theorem 4.5). This allows us to prove our main result, that the maps consisting of conjugating by a transformation matrix are precisely the trace-preserving isomorphisms of the left regular representations of $G$ and $G^{\prime}$ that commute with conjugate transpose. This implies that a $(G,G^{\prime})$-invariant quantum Latin square exists if and only if the multiset of degrees of irreducible representations of $G$ is the same as that of $G^{\prime}$, i.e., if and only if the group algebras of $G$ and $G^{\prime}$ are isomorphic. In particular, if $G$ and $G^{\prime}$ are abelian groups of the same order, such a quantum Latin square always exists and in this case there are finitely many. If, however, a $(G,G^{\prime})$-invariant quantum Latin square exists for non-abelian groups $G$ and $G^{\prime}$, then there are uncountably many of them.
@@
In Section 6 we delve into some representation theory that we will need for our main result and also investigate the relationship between transformation matrices of $(G,G^{\prime})$-invariant quantum Latin squares and what we refer to as quasi-regular representations of $G$ and $G^{\prime}$. We prove our main result in Section 7, namely that maps given by conjugating by a transformation matrix of a $(G,G^{\prime})$-invariant quantum Latin square are precisely the trace and conjugate transpose preserving isomorphisms of the group algebras of $G$ and $G^{\prime}$. Section 8 shows how to actually construct $(G,G^{\prime})$-transformation matrices, assuming one has unitaries that block-diagonalize the group algebras. In the abelian case this is achieved by the character tables of the groups, and thus we can construct all $(G,G^{\prime})$-transformation matrices since there are finitely many in this case. We study the abelian case further in Section 9 and show that in this case the convex hull of the correlations produced by $(G,G^{\prime})$-invariant quantum Latin squares is the image of the set of classical $(G,G^{\prime})$-invariant correlations under a particular unitary map. In Section 10 we show that the submatrices of $(G,G^{\prime})$-transformation matrices corresponding to the rows/columns with a single nonzero entry give isomorphisms between subgroups of $G$ and $G^{\prime}$, as well as the linear characters of these groups. Additionally, we prove a partial converse assuming a certain product structure of the groups $G$ and $G^{\prime}$. In Section 11 we define the support graph of a correlation and consider in particular support graphs of group invariant correlations. Section 12 presents some computational results and important examples, e.g., an example of a non-classical correlation produced by a $\mathbb{Z}_{2}^{4}$-invariant quantum Latin square, answering a question of [18]. Finally, we end with a discussion of our results and some open problems.
@@ 2 The Isomorphism Game
It follows from the proof of Theorem 5.4 in [2] that a quantum permutation matrix $\mathcal{P}=(p_{xy})_{x\in V(X),y\in V(Y)}$ satisfying the conditions of the above theorem results in a correlation $p$ that wins the $(X,Y)$-isomorphism game and is given by
@@
where $\operatorname{tr}$ denotes the normalized trace, i.e., $\operatorname{tr}(M)=\operatorname{Tr}(M)/d$ for $M\in M_{d}(\mathbb{C})$. It is well known that if all of the entries of a quantum permutation matrix pairwise commute, then the resulting correlation is classical. The fact that the converse is not true is the focus of [18].
In some cases, e.g. Section 11, it may be useful to allow our quantum permutation matrices to be over some self-adjoint subalgebra $\mathcal{A}\subseteq M_{n}(\mathbb{C})$. By this we mean that the entries of the quantum permutation matrix all live in $\mathcal{A}$ and the sum of any row or column is the identity of $\mathcal{A}$, which is some projection $q$ in $M_{n}(\mathbb{C})$. In this case, the normalized trace is defined as the usual trace of $M_{n}(\mathbb{C})$ divided by the trace of $q$, and the formula (2) still produces a quantum correlation.
@@
It is likely, but unknown, that not every winning quantum correlation for the $(X,Y)$-isomorphism game arises from a quantum permutation matrix via Equation (2). However, it follows from known results in the literature, see e.g. [10], that every winning quantum correlation for the $(X,Y)$-isomorphism game is a convex combination of correlations of the form (2). Importantly, the sets of classical and quantum correlations are convex with the latter containing the former. Lastly, we note that it is well-known, and easy to see from (2), that all quantum correlations satisfy the so-called non-signalling condition, i.e., that Alice’s marginals $p_{A}(x|y):=\sum_{x^{\prime}}p(x,x^{\prime}|y,y^{\prime})$ do not depend on Bob’s input $y^{\prime}$ and Bob’s marginals $p_{B}(x^{\prime}|y^{\prime}):=\sum_{x}p(x,x^{\prime}|y,y^{\prime})$ do not depend on Alice’s input $y$. This condition enforces that no information is transmitted between the two parties.
Notably, it is possible for graphs $X$ and $Y$ to be quantum isomorphic but not isomorphic [2]. In this case, the correlation of any winning quantum strategy for the $(X,Y)$-isomorphism game is necessarily non-classical. However, the converse does not hold: there exist non-classical quantum correlations that win the $(X,Y)$-isomorphism game, but only for isomorphic $X$ and $Y$, e.g., the correlation used in [18, Theorem 3.19] to show that $K_{5}$ has “nonlocal symmetry”. Note that if $p$ is a winning correlation for the $(X,Y)$-isomorphism game, then it is also a winning correlation for the isomorphism game where $X$ and $Y$ are replaced by empty graphs on the same vertex set. Since we are mainly focused on group invariant quantum Latin squares and the correlations they produce, this least restrictive isomorphism game is the one we will consider. Since for empty graphs, two vertices can only be equal or distinct non-adjacent, in this case it makes sense to simply think of the isomorphism game as a bijection game between two sets, in which the players must give the same answer if and only if they are given the same input. For us, the two sets will always be two groups $G$ and $G^{\prime}$. As with the isomorphism game, we restrict to the case where $|G|=|G^{\prime}|$, but remark that this assumption can be removed as in the isomorphism game case by passing to a larger game. It is then clear that the winning deterministic strategies for the $(G,G^{\prime})$-bijection game correspond precisely to the bijections between $G$ and $G^{\prime}$. Thus there is always a winning classical strategy, but we are interested in the winning quantum strategies that produce non-classical correlations. However, in Section 3.2 we will show how to take any winning group invariant correlation for the bijection game and determine all the graph isomorphism games for which it is also a winning strategy. Note that any quantum permutation matrix with rows indexed by $G$ and columns by $G^{\prime}$ gives a winning quantum correlation for the $(G,G^{\prime})$-bijection game, since the $A_{X}\mathcal{P}=\mathcal{P}A_{Y}$ condition of Theorem 2.4 is trivial for empty graphs.
@@ 3.2 Cayley (Di)Graphs
For any $n\times n$ doubly stochastic matrix $D$ and vectors $u,v\in\mathbb{C}^{n}$ with the same multiset of entries, it is well known that $Du=v$ if and only if $D_{ij}=0$ whenever $u_{j}\neq v_{i}$ (see, e.g., [12, Lemma 7.1]). Therefore $D^{p}\mathbf{e}_{C^{\prime}}=\mathbf{e}_{C}$ if and only if $D^{p}_{a,b}=0$ whenever ($a\in C$ and $b\notin C^{\prime}$) or ($a\notin C$ and $b\in C^{\prime}$). By definition of the graph $B$, this is equivalent to there being no edges between $C\cup C^{\prime}$ and $(G\cup G^{\prime})\setminus(C\cup C^{\prime})$. This is, in turn, equivalent to $C\cup C^{\prime}$ being the vertex set of a union of connected components of $B$, which is equivalent to $C=C_{I}$ and $C^{\prime}=C^{\prime}_{I}$ for some $I\subseteq[k]$. ∎
@@ 7 Existence of transformation matrices
By the Wedderburn-Artin Theorem [4, Theorem 1.3.5], any semisimple Artinian algebra can be written as a direct sum of full matrix algebras whose dimensions are uniquely determined. Now the result follows from Lemma 7.1. ∎
@@ 9 The abelian case
It is straightforward to check that the transformation matrix of the $(G,G^{\prime})$-invariant quantum Latin square described in Example 4.4 is precisely the matrix $\mathcal{C}^{\dagger}P^{\pi}\mathcal{C}^{\prime}$ of the above theorem, where $\pi$ is the bijection taking $\chi^{\prime}_{i}$ to $\chi_{i}$. Thus the construction described in that example, which is the generalization of the construction of [18, Definition 4.1] to the case of possibly different groups $G$ and $G^{\prime}$, produces all $(G,G^{\prime})$-invariant quantum Latin squares, up to isometry.
@@ 10 Isomorphisms
Now consider the groups $K$ and $K^{\prime}$. It is well known that the irreducible characters of a direct product of groups are precisely the Kronecker products of the irreducible characters of the factors (see for example [8, Theorem 19.18]). It follows that the multisets of degrees of the irreducible representations of $K$ and $K^{\prime}$ are the same. Therefore, there exists a $(K,K^{\prime})$-transformation matrix, $W$. We will show that that the matrix $U:=V\otimes W$ is a $(G,G^{\prime})$-transformation matrix satisfying the required conditions. Clearly, $U\in\mathbb{C}^{(H\times K)\times(H^{\prime}\times K^{\prime})}=\mathbb{C}^{G\times G^{\prime}}$, and it is a unitary matrix since $V$ and $W$ are unitary.